# Arithmetic Properties of Several Generalized-Constant Sequences, with Implications for $\boldsymbol{\Gamma^{\left(n\right)}\left(1\right)}$

### v5

@@ 5. Conclusion
$\bullet\>$ at least one element of the pair $\left\{\gamma,\gamma^{\left(2\right)}\right\}$ and at least two elements of each set $\left\{\gamma,\gamma^{\left(n\right)},\gamma^{\left(n+1\right)},\ldots,\gamma^{\left(2n\right)}\right\}$, for $n\in\left\{2,3,\ldots,26\right\}$, are transcendental, with analogous implications for the sequence $\Gamma^{\left(n\right)}\left(1\right)=\left(-1\right)^{n}\gamma^{\left(n\right)}$;
@@
$\bullet\>$ the $\delta^{\left(n\right)}/e$ are transcendental infinitely often, with the density of the set of transcendental terms having asymptotic lower bound $1/2$.
@@
$\bullet\>$ the $\widetilde{\delta}^{\left(n\right)}/e$ are transcendental infinitely often, with the density of the set of transcendental terms having asymptotic lower bound $1/2$.
@@
each of which is itself an $E$-function, and these new functions arranged into a system of linear differential equations with coefficients in $\mathbb{Q}\left(z\right)$ and a unique finite singularity at $z=0$. By employing the Kolchin-Ostrowski differential-field framework (see Srinivasan [18]), we then can show that, for any $m\in\mathbb{Z}_{\geq 1}$, the functions $G_{1}\left(z\right),H_{1}\left(z\right),G_{2}\left(z\right),H_{2}\left(z\right),\ldots,G_{m}\left(z\right),H_{m}\left(z\right)$ are algebraically independent over $\mathbb{C}\left(z,G_{0}\left(z\right),H_{0}\left(z\right)\right)$. This allows us to conclude, via the refined Siegel-Shidlovskii theorem (see Beukers [15]), that for $z=1$ the set of values $\left\{G_{1}\left(1\right),H_{1}\left(1\right),G_{2}\left(1\right),H_{2}\left(1\right),\ldots,G_{m}\left(1\right),H_{m}\left(1\right)\right\}$ are algebraically independent over $\overline{\mathbb{Q}}\left(G_{0}\left(1\right),H_{0}\left(1\right)\right)$, and therefore transcendental. Finally, recognizing that
@@
  • [4] Rivoal, T., 2012, “On the Arithmetic Nature of the Values of the Gamma Function, Euler’s Constant, and Gompertz’s Constant”, Michigan Mathematics Journal, 61, 2, 239-254.
@@
  • [7] Powers, M. R., 2025, “Conjunctions of Three ’Euler Constants’ in Poisson-Related Expressions”, arXiv:2503.02054.
@@
  • [16] Powers, M. R., 2025, “On the Uniqueness of Ein(1) among Linear Combinations of the Euler-Mascheroni and Euler-Gompertz Constants”, arXiv:2510.00315.
  • [17] Berndt, B. C. and Rankin, R. A., 1995, Ramanujan: Letters and Commentary, American Mathematical Society.
  • [18] Srinivasan, V. R., 2008, “Extensions by Antiderivatives, Exponentials of Integrals, and by Iterated Logarithms”, arXiv:0811.3004.

### v6

@@ 5. Conclusion
$\bullet\>$ the $\gamma^{\left(n\right)}$ are transcendental infinitely often and at least two elements of each set $\left\{\gamma,\gamma^{\left(n\right)},\gamma^{\left(n+1\right)},\ldots,\gamma^{\left(2n\right)}\right\}$, for $n\in\left\{2,3,\ldots,26\right\}$, are transcendental (with analogous implications for the sequence $\Gamma^{\left(n\right)}\left(1\right)=\left(-1\right)^{n}\gamma^{\left(n\right)}$);
@@
$\bullet\>$ both the $\delta^{\left(n\right)}/e$ and $\delta^{\left(n\right)}$ are transcendental infinitely often, with the densities of both sets of transcendental terms having asymptotic lower bound $1/2$.
@@
$\bullet\>$ both the $\widetilde{\delta}^{\left(n\right)}/e$ and $\widetilde{\delta}^{\left(n\right)}$ are transcendental infinitely often, with the densities of both sets of transcendental terms having asymptotic lower bound $1/2$.
@@
each of which is itself an $E$-function, and these new functions can be arranged into a system of linear differential equations with coefficients in $\mathbb{Q}\left(z\right)$ and a unique finite singularity at $z=0$. By employing the Kolchin-Ostrowski differential-field framework (see Srinivasan [18]), we then can show that, for any $m\in\mathbb{Z}_{\geq 1}$, the functions $G_{1}\left(z\right),H_{1}\left(z\right),G_{2}\left(z\right),H_{2}\left(z\right),\ldots,G_{m}\left(z\right),H_{m}\left(z\right)$ are algebraically independent over $\mathbb{C}\left(z,G_{0}\left(z\right),H_{0}\left(z\right)\right)$. This allows us to conclude, via the refined Siegel-Shidlovskii theorem (see Beukers [15]), that for $z=1$ the set of values $\left\{G_{1}\left(1\right),H_{1}\left(1\right),G_{2}\left(1\right),H_{2}\left(1\right),\ldots,G_{m}\left(1\right),H_{m}\left(1\right)\right\}$ are algebraically independent over $\overline{\mathbb{Q}}\left(G_{0}\left(1\right),H_{0}\left(1\right)\right)$, and therefore transcendental. Finally, recognizing that
@@
  • [4] Rivoal, T., 2012, “On the Arithmetic Nature of the Values of the Gamma Function, Euler’s Constant, and Gompertz’s Constant”, Michigan Mathematical Journal, 61, 2, 239-254.
@@
  • [7] Powers, M. R., 2025, “Conjunctions of Three ’Euler Constants’ in Poisson-Related Expressions”, arXiv:2503.02054v4.
@@
  • [16] Powers, M. R., 2025, “On the Uniqueness of Ein(1) among Linear Combinations of the Euler-Mascheroni and Euler-Gompertz Constants”, arXiv:2510.00315v1.
  • [17] Berndt, B. C. and Rankin, R. A., 1995, Ramanujan: Letters and Commentary, American Mathematical Society, Providence, RI, USA.
  • [18] Srinivasan, V. R., 2008, “Extensions by Antiderivatives, Exponentials of Integrals, and by Iterated Logarithms”, arXiv:0811.3004v1.